OpenStudy (anonymous):

Given: 3x + 1 = 2 + 2x – 4 Prove: x = –3

OpenStudy (anonymous):

Given the equation 3x + 1 = 2 + 2x – 4, use the commutative property to rearrange the terms so that like terms are next to one another. This gives the equation 3x + 1 = 2 – 4 + 2x. Then, use the associative property of addition to group the like terms. This gives the equation 3x + 1 = (2 – 4) + 2x. Next, combine like terms to get the equation 3x + 1 = – 2 + 2x. Use the subtraction property of equality to subtract 2x from both sides of the equation. This gives the equation x + 1 = – 2. Then use the _________________________ to subtract 1 from both sides of the equation. This gives the solution x = –3. Therefore, given the equation 3x + 1 = 2 + 2x – 4, x is equal to –3. Which justification was left out of the paragraph proof above? Answer Associative Property of Addition Commutative Property of Subtraction Addition Property of Equality Subtraction Property of Equality

OpenStudy (anonymous):

da fuq? lmao.

OpenStudy (anonymous):

lmao i know right !

OpenStudy (anonymous):

hold on ill try these

OpenStudy (anonymous):

what grade are you in? I wanna make sure i'm answering these in my grade lv. i'm a freshmaen in highschool

OpenStudy (anonymous):

im a sophomore in highschool

OpenStudy (anonymous):

ill try

OpenStudy (anonymous):

thank you !

OpenStudy (anonymous):

the first one is 1+3x=2x-4+2

OpenStudy (anonymous):

got dam i didnt read the whole question

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

hahahahahaha its okay ! this is soooo retarded

OpenStudy (anonymous):

i'm guessing the fisrt one could be the property of subtraction. I've trick questions like this before.

OpenStudy (lgbasallote):

go @trueseminole you can do it :DDDD \m/

OpenStudy (lgbasallote):

oh you did it already huh

OpenStudy (anonymous):

lmao

OpenStudy (lgbasallote):

just look at x+ 1 = -2 x = -3 what happened to the left side of the equation..

OpenStudy (anonymous):

i thought it was more than one question, but I gues it was just one long retriceone

OpenStudy (anonymous):

hahahaha ! thank you so much : )

OpenStudy (anonymous):

excuse the typos, i'm lagging

OpenStudy (anonymous):

The justification is Subtraction Property of Equality. This property says that both sides, if you subtract a common number, still equal the same thing. For example in \[1+x=2\] then if I subtract 3 to each side: \[1+x-3=2-3\] And solve the equation: \[-2+x=-1\] Notice how x is still equal to 1. In your example, you subtracted from both sides the same number, thus you haven't changed the value.